OpenStudy (anonymous):

Solve the following equation -10-3/7x=5

OpenStudy (callisto):

\[-10 -\frac{3}{7x} =5\]Is this your question?

OpenStudy (anonymous):

yeah but the x is 3/7x im solving to get x by itself

OpenStudy (anonymous):

-3/7x = 15 -3 = 105x 105x = -3 x = -3/105 = -1/35

OpenStudy (callisto):

\[-10 -\frac{3}{7}x =5\] This?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

\[-10 -\frac{3}{7}x =5\]\[ -\frac{3}{7}x =5+10\]\[ -\frac{3}{7}x =15\]\[ -\frac{3}{7}x (\frac{-7}{3})=15(\frac{-7}{3})\]\[ x=5(-7) = -35\]

OpenStudy (anonymous):

thanks